40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 16 01:03:59 2011, Jenne, Update, IOO, MC alignment work 
    Reply  Thu Feb 17 14:52:27 2011, Jenne, Update, IOO, MC alignment work 
Message ID: 4316     Entry time: Thu Feb 17 14:52:27 2011     In reply to: 4305
Author: Jenne 
Type: Update 
Category: IOO 
Subject: MC alignment work 

I worked a little bit more on optimizing the mode matching to the MC, but it's still not great.  I've only gotten a visibility of ~45%, but Koji said that it used to be ~87%.  So there is a long way to go.  Kiwamu said he can work with the lower-power configuration for a few days, and so my next step will be to measure the beam profile (stick a window in the path, and look at the refl from the window....that way we don't get thermal lensing from transmission through an optic), and redo the mode matching calculation, to figure out where the last lens should actually sit.

Quote:

So.... Kiwamu and I were concerned (still a little concerned) that ETMY is not damping as nicely as it should be.  (It's fine, but the UL rms is ~5, rather than ~1 or less. BURT restores by Kiwamu didn't change anything.) Anyhow, I was heading out to push the annoying ribbon cables more firmly into the satellite adapter board things that are tied to the racks in various places (The back of 1X5 for the corner optics and the end station racks for the ETMs).  The point was to push in the ETMY one, but while I was out in the lab and thinking about it, I also gave all of the corner connectors (MC1, MC2, MC3, ITMx, ITMY, BS, PRM, SRM) a firm push. 

Kiwamu noticed that when I did this, the Mode Cleaner alignment got a little bit worse, as if the connection to the satellite adapter boards hadn't been great, I pushed the connectors in and the connection got better, but we also got a bit of a DC offset in the MC alignment.  Anyhow, the MC_TRANS power went down by ~2, to about the place it had been before Kiwamu adjusted the position of the lens in between the zigzag mirrors.  (I don't know if Kiwamu elogged it earlier, but he scooted the lens a teensy bit closer in the optical path to the Mode Cleaner). 

To counteract this loss in MC transmitted power as a result of my connector actions, I went back to the PSL table and fiddled with the zigzag steering mirrors that steer the beam from the PSL table over to the mode cleaner.  I got it a little better, but it's still not perfect.

Kiwamu has noted that to improve the mode matching into the Mode Cleaner with the new PMC in place, we might have to move the lens which is currently between the zigzag steering mirrors, and put it after the second mirror (so in between the last steering mirror and the pickoff window that sends a piece of the beam over to PSL_POS and PSL_ANG).  This will make the waist between MC1 and MC3 tighter. 

Moral of the story:  To improve IMC mode matching we need to move the last lens closer in the optical path to the mode cleaner waist. Twiddle with zigzag steering mirrors to optimize.

 

ELOG V3.1.3-